solve this equation 5x+9=3x+41
x=

Answer by lenny460(1073)

solve this equation 5x+9=3x+41
x=

5x + 9 = 3x + 41
Subtract 3x from each side of the equation.
5x + 9 - 3x = 3x + 41 - 3x
Combine like terms:
5x - 3x = 2x
3x - 3x = 0
Therefore:
2x + 9 = 41
Subtract 9 from each side of the equation.
2x + 9 - 9 = 41 - 9
Combine like terms:
9 - 9 = 0
41 - 9 = 32
2x = 32
Divide each side by 2
2x/2 = 32/2
x = 16

Check:
5x + 9 = 3x + 41
5(16) + 9 = 3(16) + 41
5(16) = 80
3(16) = 48
80 + 9 = 89
48 + 41 = 89
89 = 89


Lennox Obuong
Algebra Tutor